OpenStudy (anonymous):

Solve the system, using substitution. Write the solution as an ordered pair leaving no spaces or write no solution or infinitely many solutions Example: (2,1) 4x + y = 4 and y = 4 - 4x

OpenStudy (anonymous):

since y=4-4x, you can substitute it into the first equation 4x+(4-4x)=4 4x+4-4x=4 so 0=0. Therefore there are infinitely many solutions. If you draw the graphs for these equations you will see that they are on top of each other.

OpenStudy (anonymous):

Ok : )